# RASP: A Drone-based Reconfigurable Actuation and Sensing Platform for Engaging Physical Environments with Foundation Models

### v2

@@ 3. Modular and Reconfigurable Sensing and Actuation Drone
Figure 2. RASP architecture.
Figure 2 shows the architecture of RASP. At a high level there are three major conceptual architectural components: the ground station or mechanical layer, electrical connection layer, and the software layer. These conceptual architectural components are implemented across five physical platform components: the ground station (Figures 3e and 4), sensor/actuation modules (Figures 3c-d), unified module interface (UMI) (Figure 5), the carrier board (Figure 3b), and the physical drone (Figure 3a).
3.1. Architecture
Mechanical Layer. On a functional level, the mechanical layer separates RASP from existing reconfigurable drone platforms by enabling the autonomous swapping of sensor and actuator modules. The drone lands on the ground station, which will physically detach the current module and attach the newly requested one. We propose novel swapping, including an augmented funnel shaped landing structure, to enable autonomous human-free swapping of modules as discussed in Section 3.2.
Electrical Connection Layer. This layer maintains the data, communication, and power connection between each sensor or actuation module with the drone, while being both detachable yet reliable mid mission. Existing reconfigurable drone platforms (da Silva Ferreira et al., 2020) generally incorporate slotted mechanical mechanisms. While these pins can maintain connection under great amounts of force, they are difficult to attach and remove, requiring a person to manually connect or disconnect. Attaching often requires precisely aligning slots or pins and applying acute pressure at junction points.
Instead, we leverage magnets to maintain connection. Attaching a magnet-based connector only requires moving the module close to the drone’s connector before the magnetic pull aligns all connections in place. Removing the module can be achieved without needing to apply acute force directly at the junction. Both of these characteristics allow magnet-based connectors to be more conducive for autonomous reconfigurable systems than purely mechanical connectors. We discuss in more detail in Section 3.2.
Software Layer. We propose standardized protocols and software libraries on top of the mechanical and electrical connection layers that allow the drone to discover, control, and receive data from attached modules on-the-fly.
3.2. Platform Components
@@
Drone platform. We build RASP off the open-source Crazyflie drone (cra, 2023), as shown in Figure 3. We chose the lightweight (27g) Crazyflie as our foundation due to its compact size, making it easier to operate in all environments, especially indoors. To increase its payload, we replaced its brushed motors (19000 Kv powered by a 1-cell battery) with more powerful brushless motors (3800 Kv powered by a 4-cell 850mAh battery). We also printed a protective guard to protect its blades from objects and people. The drone can hover for nearly 4 minutes, which is comparable to the Crazyflie.
Carrier board. The carrier board, which is attached to the base of the drone, provides the physical data, power, and communication connections between the drone platform and each sensor/actuation module. We implement the carrier board on a lightweight \$15 Raspberry Pi Zero 2W (rpi, 2023), which has connections for one sensor or actuation module.
@@
Table 1. RASP’s supported sensors/actuators and their mass, power consumption, and cost.
@@
Sensor, Actuation, and Recharging Modules. RASP comes with a collection of sensor and actuation modules (full list in Table 1). These modules are attached or detached from the drone by the ground station/mechanical layer and comes with a 3D printed structure that increases the surface for the ground station to pick up modules. The actuation module, shown in Figure 3d, is a container structure with a motor that opens and closes a hatch. Small items (e.g., medication, candy, pet food, etc.) can be loaded into this module for the drone to deliver.
Additionally, we create a module that recharges the battery of the drone, which gets attached and removed just like other sensors and actuators. Unlike other standalone sensors and actuators, the recharging module is connected to the wall-powered ground station and charges the drone’s battery at a rate of 1.7 A. At this rate, the fully depleted drone can be fully charged in 30 minutes. The recharging module is the default module that’s attached to the drone each time the drone finishes a task, returns to the ground station, and has no other pending tasks.
@@
Ground station and automated takeoff and landing. The ground station primarily spans the mechanical layer, consisting of a gripper mechanism and a conveyor belt to position modules to swap in/out. We created the platform leveraging the chassis of the open-source Ender-3 3D printer (end, 2023) because it provides enough space for the drone to land, as well as enough mechanical control to remove onboard sensors and attach new sensors onto the drone. The gripper in Figure 3e removes and attaches modules onto the drone.
The ground station contains two additional components: the communication hub and the landing station. The communication hub is the central processor that controls the swapping process. Drones communicate with this hub prior to landing to specify which module it requires and to initiate the swapping process.
@@
The drone then lands at the landing station before the mechanical layer begins the swapping process. The communication hub guides the drone into position for landing using two cameras facing up on top of the ground station (Figure 4b). We also print and attach two ArUco markers to the bottom side of the drone to easily detect the position of the drone and its orientation. ArUco markers, commonly used for camera pose estimation, are similar to QR codes, but carry less encoded information, which makes them more computationally efficient to detect (Sani and Karimian, 2017). In future work, we plan to leverage more complex computer vision models to automatically determine the position and orientation of the drone without needing to add additional markings.
Landing Platform. Removing an attached sensor and ensuring that the new sensor is securely fastened requires the drone to land precisely at the specific location where the interchange occurs. However, precise drone landing and alignment is extremely difficult because nonlinearities in drone stability become more pronounced as the drone moves closer to the ground (ground effect) (Matus-Vargas et al., 2021). Rather than creating a flat landing pad that is similar to the landing pads for helicopters, we take inspiration from the Ring Always Home Drone, which uses a funnel (rin, 2023). As long as the drone lands within the larger radius of opening of the funnel, the drone will automatically slide towards the bottom of the funnel with an opening that latches onto the onboard module (Figure 4c-top). To further improve alignment, we propose a new design that includes additional grooves to further improve module and drone alignment (Figure 4c-bottom).
We benchmark both the vanilla funnel and the grooved funnel landing platform design, as shown in Figure 4d. For each design, we had the drone land and swap modules 75 times and found that the swap rates are fairly similar. However, the grooved design corrects for greater drone misalignments from the platform, as reflected in the higher maximum tolerable orientation and offset errors. As such, we adopt our proposed grooved funnel design into RASP.
Figure 5. Universal module interface for connecting and disconnecting sensors and actuators, as well as recharging the drone.
Unified Module Interface (UMI). The UMI (Figure 5) spans both the electrical connection layer and the software layer, giving the carrier board on the drone a standardized way to communicate with a variety of sensing and actuation modules. The UMI has the following design features.
Standardized 24-pin magnetic connections: Unlike existing reconfigurable drone and sensing platforms that use mechanical connectors, each sensor and actuation module in RASP comes equipped with two magnetic connectors each with 12 spring-loaded pins, with corresponding connectors on the carrier board. The magnets allow each module to remain attached and connected to the drone during flight, while being pliable enough to be easily removed by the mechanical layer. While mechanical pins and slots are the standard for maintaining stable connections, they are difficult to attach and remove without human intervention. Magnets allow attaching and removing modules with much less precision, with enough force to maintain connection mid-mission. Magnetic connections only require placing the module in the vicinity of the drone to attach the module and diffused force to remove, compared to standard mechanical connectors.
Sensor and actuator drivers: To support a variety of sensors and actuators that may each be interfaced differently, RASP implements a library of standard drivers, as part of the software layer, that support different communication and sampling schemes, namely: I2C, SPI, UART, GPIO, and ADC. We also provide developers with a driver template that allows them to easily incorporate new sensors and actuators into the RASP ecosystem.
Plug-and-play controller: Once a new module attaches to the carrier board, the drone needs to be able to start collecting data or using the actuator automatically, without needing to manually run any command. To enable this feature, we include a non-volatile memory chip on each module (M24C02 (STMicroelectronics, 2017) 2Kb I2C bus electrically erasable programmable read-only memory (EEPROM)), which stores the identity of the module. The carrier board will pull data from here every time a new module is attached and automatically downloads the corresponding driver.
4. Towards a Fully Autonomous Drone-based Personal Assistant.
Figure 6. System architecture of intelligent assistant with RASP.
In this section, we detail the prototype of RASP into a personal assistant application that has information about the space through a deployment of static sensors (namely cameras). Sections 4.1 and 4.2 introduce methods for adapting visual language models to analyze and understand sensor data from the environment, while Sections 4.4 and 4.3 introduce how FMs can integrate RASP into their pipeline.
@@
Figure 6 shows the workflow of our prototype. Static cameras in the environment provide a high-level and coarse-grained view of the environment; in future work and systems, we envision RASP being compatible with other static sensors, beyond cameras. When a user gives a voice command, the system leverages a visual-language model to analyze the scene, interpret the command, and output a list of potential locations of interest, along with the relevant sensor or actuator. Then, the system actuates the RASP drone with the sensor/actuator to each of the locations to complete the task.
For example, one task or question a person may want to know is “where is the warmest place to sit”? A standard RGB camera would not be able to accurately make this determination without additional sensors (e.g., temperature). The voice command and images are analyzed by a visual-language model, which outputs potential locations where a person could sit (e.g., chairs and couches). Then, the system outfits the RASP drone with a temperature module and actuates the drone to each of these locations to measure temperature. The time series data from the temperature sensor will then be analyzed by another language/foundational model to make the final determination. There are four classes of tasks that we explore in this deployment, multiple of which are currently not commonly packaged into existing IoT smart devices:
Object/Location Identification. This set of tasks requires the system to observe the environment and identify an object or location based on the user’s command. For example, the command “where is my phone” requires the visual-language model to identify potential locations for where the phone could be, before sending a drone with a camera to take a closer look and verify.
Object/Location State. This set of tasks involves learning about the state or condition of a specific object or location. For example, “is my food burning” would require the system to identify food that is being cooked, before sending a drone with a CO2 sensor to determine the state of the food.
Surveillance. These tasks differ from the previous two, as it requires the visual-language model to continuously analyze the scene until a potential event occurs, before sending the drone (e.g., “Let me know if any chemical spills on the table”).
Actuation. Unlike the previous three categories, all of which require the drone to sense the environment, actuation tasks all leverage the actuation module to bring an item (e.g., medicine, candy, food) to a specific location.
4.1. Setup and Networked Camera System
To effectively use a camera network to observe the environment, there generally is a lengthy setup process where you need to specify the location of the cameras and have the floor map of the building to stitch together different views. We adopt this self localizing camera network (Liu et al., 2022) that allows users to semi-randomly place cameras on the ceiling, facing downwards perpendicular to the floor; by leveraging motion and people walking throughout daily life, the camera network automatically self-localizes each camera within the network and generates the walkable floormap of the areas observed by the cameras. In this way, users only need to install cameras and do not need to define a floormap or specify and calibrate based on location. More details of our deployment are in Section 5.
4.2. Visual-Language Models for Interpreting Voice Commands
Figure 7. Summary of prompts used to engineer the LLaVA visual-language model (left) to interpret and identify locations of interest given a voice command from a user, as well as for ChatGPT (right) to analyze time series data from the RASP onboard drone sensor.
Figure 8. Segmentation and clustering to break down scenes into smaller more manageable pieces for LLaVA and DINO. (a) Object masks after applying Segment Anything Model (SAM). Extracted frames after clustering object masks based on K-Means (b), hierarchical clustering (c), and ARCK-Means (d). For ARCK-Means, we constrain the aspect ratio of extracted frames to be between 0.67 and 1.5.
Identifying Objects of Interest. When a user issues a voice command, we pass this command along with the full image of the environment stitched together from all cameras into the open-source Large Language-and-Vision Assistant (LLaVA) visual-language model (Liu et al., 2023a). For the surveillance category, we pass in two consecutive frames or segments to detect changing events. LLaVA outputs (in text) the names of potential objects in the scene that might of interest, as well as the suitable sensor/actuation module that the drone should use. Figure 7-left shows the series of engineered and task-specific prompts we leverage, which was designed based off of the highly active Brex LLM prompt engineering guide (Brex Inc., 2023). The LLaVA text outputs are input as a prompt into Grounding DINO (Liu et al., 2023b), a state-of-art language model for zero-shot object detection, which then outputs bounding boxes and locations of the specified objects.
Object Segmentation and Clustering. However, we found that LLaVA was not able to reliably detect objects of interest in a scene when we passed in an image or a large space (e.g., a single chair in a large room); the larger the scene, the less detailed and more high-level the responses became. To capture small details within the scene that might be relevant to a voice command, we used the Segment Anything Model (SAM) to extract key objects and smaller areas of the scene (Kirillov et al., 2023). Figure 8a shows an example of the object masks that are output by SAM.
We tried inputting these masked objects into LLaVA and then DINO one-by-one, but there were two issues we noticed. First, many of the masks generated often splits objects between two different masks; as such, DINO would often detect the same object multiple times. Second, the masks are not shaped like a regular image (e.g., a square or rectangle); we saw a large degradation in detection performance if we zero-padded the areas around the masked object into a rectangular shape.
To rectify this, we cluster multiple segments together to ensure potentially important objects are all contained with a single rectangular segment. We benchmark the following clustering methods: K-means (Ahmed et al., 2020), hierarchical clustering (Murtagh and Contreras, 2012), and Aspect Ratio Constrained K-Means (ARCK-Means). For K-means and hierarchical clustering, we cluster the centroids of each mask from SAM into $k$ different clusters. Next, we find the minimum-area rectangle that encompasses all masks of each cluster. These $k$ minimum-area rectangles are then analyzed by LLaVA and DINO.
One of the challenges we found with DINO is that it was not able to identify objects of interest very well if input images are very long or very wide. As such, we create and benchmarked a new clustering method called Aspect Ratio Constrained K-Means (ARCK-Means), which constrains the aspect ratio of the clustered masked between two values $x<y$ to maintain a more square shape. To accomplish this, we check if adding a new segment into the existing cluster (during hierarchical clustering) causes the aspect ratio to fall below or above our constraints (e.g., $<x$ or $>y$); If these constraints are broken, then we do not make the assignment.
First, we tested two prompts into the LLaVA model and vary the size of the object with respect to the frame of one camera (Figure 9a). The first prompt is more general, asking LLaVA to describe the objects present. The second prompt is more specific, asking if the specific object of interest is present in the scene. We see that the more specific a prompt is, the higher the recall, which is the reason why we used more specific prompting (Figure 7). Second, we see that as the size of the object gets smaller with respect to the image frame, the recall gets smaller, since the signal-to-noise ratio of the object gets smaller.
In Figure 9b, we compare the recall of each clustering method after clustering each scene into 5 segments, and see that ARCK-Means has the highest recall due to improvements in the masks it clusters and the shape of the resulting segments; as such, we adopt ARCK-Means into the final system. Figure 9c shows the run time and recall of ARCK-Means as a function of the number of clusters or segments we split the scene into. We see that the recall levels out at around $90\%$ after $5$ clusters, while taking around $100$ms to run the full pipeline. Adding in more clusters does not yield significant improvements in detection, but significantly increases run time; as such, we segment the scene into five segments in the final system. To generate these plots, we used 47 images of indoor home, office, and lab environments. We took 35 images from the ADE20K scene parsing dataset (Zhou et al., 2017) as well as 12 images from our own deployment. We implement and run the full visual-language model pipeline (SAM, clustering, LLaVA, and DINO) on an Nvidia RTX 3090 GPU server.
@@ 4.2. Visual-Language Models for Interpreting Voice Commands
4.3. Drone Actuation and Analysis
Once the visual-language models determine important locations, the system attaches the relevant sensor or actuator (from LLaVA) to the RASP drone before moving the drone to each of these locations. In the case of an actuation task, the drone flies to the specified location, opens the hatch to drop the item of interest, before flying back to the ground station.
In the case of a sensing task, the drone flies to each location while reading sensor values from the attached module. After the drone lands back at the ground station, the time series data is input into the GPT-4 language model (OpenAI, 2023) to analyze and make the final determination. Figure 7-right shows the prompts we used for GPT-4. If the sensing modality is to use a camera to identify an object (e.g., a objection/location ID task); the image from the drone is input back into the Grounding DINO model to make the final determination.
@@ 4.4. Drone Navigation
We use images from the camera network to guide the drone by attaching an aRuco marker to the top of the drone, just like the patterns used for landing the drone in Section 3. To move the drone to specific locations, we use the straight line path from the drone’s current location to the closest point of interest. Throughout our deployments in Section 5, we observed an median 2-D localization error of 3.29cm, using our camera network. Moreover, as shown in Figure 10, the localization error does not increase over time, as is common in inertial measurement unit and dead reckoning approaches. While there are more practical approaches for drone navigation, the focus of this work is to demonstrate the utility of RASP to LLMs that interact with the physical world. Hence, we leave these aspects to future work.
5. Deployment
We deployed our RASP based personal assistant into an office/lab setting as shown in Figure 10e, using a network of 4 cameras. As discussed in Section 4.1, the cameras self-localized themselves and generated the floor map according to (Liu et al., 2022) after an afternoon of observing people moving through the space. The goal of this deployment is to demonstrate improvements in task completion rate RASP provides to FMs for more general and less structured applications.
@@ 5.1. Benchmarking
Refer to caption Refer to caption
Figure 12. Breakdown of accuracy and false positive detections by sensing task category with and without RASP. We see that leveraging RASP in conjunction with static cameras greatly reduces false detections and improves accuracy because the drone can get a closeup view or sense an important part of the environment that a camera alone cannot (e.g., humidity).
Camera Only Camera + RASP
Scenario Precision Recall F-1 Accuracy Sensor Used Precision Recall F-1 Accuracy
@@
Find Phone 68.85% 84.00% 75.68% 73.00% Drone Cam 100.00% 84.00% 91.30% 92.00%
Find Key 78.05% 80.00% 79.01% 71.67% Drone Cam 100.00% 80.00% 88.89% 86.67%
Sit - Temperature 23.47% 92.00% 37.40% 25.24% Temperature 76.67% 92.00% 83.64% 91.26%
Sit - Humidity 25.81% 88.89% 40.00% 28.00% Humidity 82.76% 88.89% 85.71% 92.00%
Sit - Light 20.62% 83.33% 33.06% 22.12% Light Sensor 95.24% 83.33% 88.89% 95.19%
Average (ID) 43.36% 85.64% 53.03% 44.01% 90.93% 85.64% 87.69% 91.42%
@@
Faucet Open 80.00% 97.78% 88.00% 82.86% Humidity 93.62% 97.78% 95.65% 94.29%
Stove Open 79.22% 87.14% 82.99% 77.27% Temperature 96.83% 87.14% 91.73% 90.00%
Average (State) 79.61% 92.46% 85.50% 80.06% 95.22% 92.46% 93.69% 92.14%
@@
Food Burning 50.91% 70.00% 58.95% 51.25% PM 93.33% 70.00% 80.00% 82.50%
Chemical Spill 25.40% 80.00% 38.55% 43.33% Gas (Alcohol) 88.89% 80.00% 84.21% 93.33%
Average (Sur.) 38.15% 75.00% 48.75% 47.29% 91.11% 75.00% 82.11% 87.91%
Average (all) 46.54% 83.17% 55.70% 48.99% 91.93% 84.79% 87.78% 90.80%
Table 2. Summary of end-to-end performance between RASP and camera-only for all sensing tasks.
@@
6) Food. In this surveillance task, users ask “let me know when my food is burning”. For safety, we simulate burning food by boiling water in a kettle, releasing water vapors. The system will then actuate the drone attached with a particulate matter sensor.
@@
For non-actuation tasks, Figure 11 shows a breakdown of the visual-language model performance in identifying the locations to send the drone versus the segmentation clustering method (Section 4). We see that the ARCK-Means clustering method yields the highest recall, across all scenarios, due to improvements in maintaining a good aspect ratio and object grouping over other clustering methods.
Figure 12 highlights improvements that RASP brings compared to a purely static camera-based system across all sensing tasks, especially with respect to false positives. A successful or “accurate” trial in this context means that the system was able to correctly identify the correct object or location (object/location ID task), correctly identify the state of the object or location (object/location state task), or correctly identify when a targeted event occurs (surveillance task); any additional points of interest identified are counted as incorrect identifications (false positives). For the non-RASP solution, we run the entire SAM, LLaVA, DINO pipeline, but do not actuate the drone. Because we are often looking for small item(s) and locations in a large scene, the visual-language model pipeline often identifies multiple points of interest (e.g., DINO draws multiple bounding boxes and locations). Without a platform such as a drone that can “zoom in” and confirm, the sensing capabilities of this system is limited, and the false positive rate becomes extremely high and the precision becomes low with additional locations identified. However, adding in the RASP-equipped drone allows the system to actuate the drone to each location to obtain a closeup view of the location and remove extraneous locations or sense an aspect of the environment that a camera cannot (e.g., humidity). This both reduces false positives and improves overall accuracy. Table 2 breaks down the recall, precision, and f-1 score across all individual tasks to further illustrate improvements in true detection rate (recall). In total, integrating RASP with FMs improved the task success rate by $85\%$.

### v3

@@ 5. Modular and Reconfigurable Sensing and Actuation Drone
In this section, we introduce the reconfigurable and modular sensing and actuation platform for drones, which FMs can actuate to identified locations of interest (Section 4).
@@
5.1. Challenges and Design
The primary challenges in realizing an automated and reconfigurable sensing platform for drones is two fold.
1. Drone and module connector. To allow reconfiguration, the sensor and actuation modules need to be detachable from the drone’s main body. While there are drone platforms that can be reconfigured with different frames, wings, or motors (Derrouaoui et al., 2022; Schiano et al., 2022; da Silva Ferreira et al., 2020), the vast majority leverage mechanical connectors with locking mechanisms that often requires additional force and complex mechanisms to fasten and remove (e.g., grippers with joints much like human hands). Instead, we leverage a fully magnetic connector. The advantage of this design choice is two-fold.
First, it simplifies the design of the mechanisms for removing and fastening new modules on the mechanical layer. Second, less force is required to fasten and remove the module. Compared to a mechanical design, such as (da Silva Ferreira et al., 2020), that requires the application of force with several mechanisms at multiple locations, potentially damaging the module, connector, or drone, a magnetic connector enables gripper to bring a new module within vicinity of the drone before the magnet automatically aligns the pins and fastens the module in place. Removing the module follows a similar procedure. We tested the success rate in swapping a module on a drone with both magnetic (Amazon, 2024) and standard PCB mezzanine connectors (Amphenol ICC, 2020a, b) using our gripper (Figure 6e) that actuates vertically to bring modules up and down. We ran 10 trials and found that our system could swap in new modules successfully every time when the connectors are magnetic. However, our system failed at swapping modules with the standard mezzanine connectors because our mechanical grippers could not apply enough force to remove nor attach modules.
2. Alignment. Swapping in new modules onto the drone can only occur successfully if the drone is oriented correctly with the gripper after landing. This challenge is not directly related to the well studied problem of drone landing (Wang et al., 2022), which aims to guide the drone to the landing platform before descending. As the drone descends closer to the ground, nonlinearities in drone stability (ground effect) (Matus-Vargas et al., 2021), for which there is no promising method to compensate, can cause the drone to become misaligned. Instead, we focus on the design of the landing platform that automatically calibrates the orientation of the drone as it touches down.
Rather than creating a flat landing pad that is similar to the landing pads for helicopters, we take inspiration from the Ring Always Home Drone, which uses a funnel (rin, 2023). As long as the drone lands within the opening of the funnel, the drone will automatically slide towards the bottom of the funnel with an opening that latches onto the onboard module (Figure 7c-top). To further improve alignment, we propose a new design that includes additional grooves to further improve module and drone alignment (Figure 7c-bottom).
We benchmark both the vanilla funnel and the grooved funnel landing platform design, as shown in Figure 7d. For each design, we had the drone land and swap modules 75 times and found that the swap rates are fairly similar. However, the grooved design corrects for greater drone misalignments from the platform, as reflected in the higher maximum tolerable orientation and offset errors. As such, we adopt our proposed grooved funnel design into FlexiFly.
5.2. Platform Components and Implementation Details
This section discusses the implementation of the rest of the reconfigurable platform.
Drone platform. We build FlexiFly off the open-source Crazyflie drone (cra, 2023), as shown in Figure 6. We replaced its brushed motors with more powerful brushless motors (3800 Kv powered by a 4-cell 850mAh battery) . To improve payload capacity and ensure operational safety in areas with human presence, we designed and fabricated a enclosed drone frame using lightweight foaming polylactic acid (LW-PLA), as shown in Figure 6a.
Carrier board. The carrier board, attached to the base of the drone, provides the physical data, power, and communication connections between the drone platform and each sensor/actuation module. We implement the carrier board on a lightweight \$15 Raspberry Pi Zero 2W (rpi, 2023), which has magnetic connections for one sensor or actuation module.
@@ 5.2. Platform Components and Implementation Details
Table 1. FlexiFly’s supported sensors/actuators and their mass, power consumption, and cost.
@@
Sensor and Actuation Modules. FlexiFly comes with a collection of sensor and actuation modules (full list in Table 1). These modules are attached or detached from the drone by the ground station and comes with a 3D printed structure that increases the surface for the ground station to pick up modules. The actuation module, shown in Figure 6d, is a container structure with a motor that opens and closes a hatch. Small items (e.g., medication, candy, pet food, etc.) can be loaded into this module for the drone to deliver.
@@
Ground station and automated takeoff and landing. The ground station consists of a gripper mechanism and a conveyor belt to position modules below the drone. We created the platform leveraging the chassis of the open-source Ender-3 3D printer (end, 2023). The gripper in Figure 6e removes and attaches modules onto the drone.
@@
To land, the ground station guides the drone using two cameras facing up on top (Figure 7b). We also print and attach two ArUco markers to the bottom side of the drone to easily detect the position of the drone and its orientation. ArUco markers, commonly used for camera pose estimation, are similar to QR codes, but carry less encoded information, which makes them more computationally efficient to detect (Sani and Karimian, 2017). In future work, we plan to leverage more complex computer vision models to automatically determine the position and orientation of the drone without needing to add additional markings.
6. Implementation of a Foundation Model and Drone-based Assistant
We show how FMs can leverage our novel methods for identifying local areas of interest (Section 4) and reconfigurable drone platforms (Section 5) to better understand and interact with physical environments, with the implementation and evaluation of a personal assistant system.
@@ 6. Implementation of a Foundation Model and Drone-based Assistant
Figure 2 shows the workflow of our prototype. Static cameras in the environment provide a high-level and coarse-grained view of the environment, just like in our preliminary deployment (Section 3.1). When a user gives a voice command, the system leverages LLaVA and ARCK-Means segmentation (Section 4) to identify local areas of interest and the relevant sensor or actuator. Then, the system actuates the FlexiFly drone with the sensor/actuator to each location to complete the task. We carry out and evaluate this system on the four classes of tasks we identify in Section 3.2.
Implementation Details. A local server runs the Ollama framework, hosts the LLMs (Llama-3.1-8B), VLM (LLaVA 1.6-8b), and Grounding DINO open-set object detection model. Upon user request, a snapshot is taken from the ceiling camera network and stitched into a single image. The server identifies required sensing modules and key areas of interest, manages pipeline execution, and sends commands to the drone. A validation LLM (Llama-3.1-8B) is used that ensures properly formatted outputs throughout the process. For sensing tasks (ID or State tasks from Section 3.2), the drone flies to each location while reading sensor values from the attached module. If the sensing modality involves camera-based object identification (e.g., an object or location ID task), the captured drone image is processed on our local server. However, for time-series data analysis, the system uses GPT-4 (OpenAI, 2023), as we found that locally hosted LLMs often produce unreliable results when generating code to analyze the sensor data. In an actuation task, the system attaches the actuation module with the relevant payload (e.g., snack or medicine) before flying to the destination for dropoff.
@@ 6.1. Drone Navigation
During flight, we use images from the camera network to guide the drone by attaching an ArUco marker to the top of the drone, just like the patterns used for landing the drone in Section 5. To move the drone to specific locations, we use the straight line path from the drone’s current location to the closest point of interest. Throughout our deployments in Section 7, we observed an median 2-D localization error of 3.29cm, using our camera network. Moreover, as shown in Figure 8, the localization error does not increase over time, as is common in inertial measurement unit and dead reckoning approaches. While there are more practical approaches for drone navigation, the focus of this work is to demonstrate the utility of FlexiFly to LLMs that interact with the physical world. Hence, we leave these aspects for future work.
7. Deployment and Evaluation
We deployed our system into an office/lab setting as shown in Figure 8e, just as in our preliminary study. The goal of this deployment is to demonstrate improvements in task completion rate FlexiFly provides to FMs for more general and less structured applications.
@@ 7.1. Benchmarking
Refer to caption Refer to caption Refer to caption
Figure 10. Breakdown of accuracy and false positive detections by sensing task category with and without FlexiFly. We see that leveraging FlexiFly in conjunction with static cameras greatly reduces false detections and improves accuracy because the drone can get a closeup view or sense an important part of the environment that a camera alone cannot (e.g., humidity).
Camera Baseline Camera + ARCK-Means only (Section 4) Camera + FlexiFly (Sections 4 and 5)
Scenario Precision Recall F-1 Accuracy Precision Recall F-1 Accuracy Sensor Used Precision Recall F-1 Accuracy
@@
Find Phone 33.33% 26.00% 29.21% 37.00% 68.85% 84.00% 75.68% 73.00% Drone Cam 100.00% 84.00% 91.30% 92.00%
Find Key 26.67% 10.00% 14.55% 21.67% 78.05% 80.00% 79.01% 71.67% Drone Cam 100.00% 80.00% 88.89% 86.67%
Sit - Temperature 15.91% 56.00% 24.78% 17.48% 23.47% 92.00% 37.40% 25.24% Temperature 76.67% 92.00% 83.64% 91.26%
Sit - Humidity 20.45% 66.67% 31.30% 21.00% 25.81% 88.89% 40.00% 28.00% Humidity 82.76% 88.89% 85.71% 92.00%
Sit - Light 20.88% 79.17% 33.04% 25.96% 20.62% 83.33% 33.06% 22.12% Light Sensor 95.24% 83.33% 88.89% 95.19%
Average (ID) 23.45% 47.57% 26.58% 24.62% 43.36% 85.64% 53.03% 44.01% 90.93% 85.64% 87.69% 91.42%
@@
Faucet Open 72.83% 74.44% 73.63% 65.71% 80.00% 97.78% 88.00% 82.86% Humidity 93.62% 97.78% 95.65% 94.29%
Stove Open 69.49% 58.57% 63.57% 57.27% 79.22% 87.14% 82.99% 77.27% Temperature 96.83% 87.14% 91.73% 90.00%
Average (State) 71.16% 66.51% 68.60% 61.49% 79.61% 92.46% 85.50% 80.06% 95.22% 92.46% 93.69% 92.14%
@@
Food Burning 44.64% 62.50% 52.08% 42.50% 50.91% 70.00% 58.95% 51.25% PM 93.33% 70.00% 80.00% 82.50%
Chemical Spill 18.03% 55.00% 27.16% 34.44% 25.40% 80.00% 38.55% 43.33% Gas (Alcohol) 88.89% 80.00% 84.21% 93.33%
Average (Sur.) 31.34% 58.75% 39.62% 38.47% 38.15% 75.00% 48.75% 47.29% 91.11% 75.00% 82.11% 87.91%
Average (all) 31.18% 51.74% 34.46% 32.17% 46.54% 83.17% 55.70% 48.99% 91.93% 84.79% 87.78% 90.80%
Table 2. Summary of end-to-end performance between FlexiFly and camera-only for all sensing tasks.
@@
6) Food. In this surveillance task, users ask “let me know when my food is burning”. We simulate burning food by boiling water. The system will then actuate the drone attached with a particulate matter sensor.
@@
7.2. Results and Analysis
We compare FlexiFly against two baselines: camera + LLaVA setup from our preliminary study (Camera Baseline) and augmenting the camera baseline with ARCK-Means only (Camera + ARCK-Means). For non-actuation tasks, Figure 9 shows a breakdown of the visual-language model performance in identifying the locations to send the drone versus the clustering method (Section 4). We see that the ARCK-Means clustering method yields the highest recall, across all scenarios, due to improvements in maintaining common aspect ratios and whole objects over other methods.
Figure 10 highlights improvements that FlexiFly brings compared to a purely static camera-based system across all sensing tasks. A successful or “accurate” trial in this context means that the system was able to correctly identify the correct object or location (object/location ID task), correctly identify the state of the object or location (object/location state task), or correctly identify when a targeted event occurs (surveillance task); any additional points of interest identified are counted as incorrect identifications (false positives). Because we are often looking for small item(s) and locations in a large scene, the visual-language model pipeline often identifies multiple points of interest (e.g., DINO draws multiple bounding boxes and locations). Without a platform such as a drone that can “zoom in” and confirm, the sensing capabilities of this system is limited, and the false positive rate becomes extremely high, and the precision becomes low with additional locations identified. However, adding in the FlexiFly-equipped drone allows the system to actuate the drone to each location to obtain a closeup view of the location and remove extraneous locations or sense an aspect of the environment that a camera cannot (e.g., humidity). This both reduces false positives and improves overall accuracy. Table 2 breaks down the recall, precision, and f-1 score across all individual tasks to further illustrate improvements in true detection rate (recall). In total, integrating FlexiFly with FMs improved the task success rate by $85\%$.